OpenStudy (anonymous):

Solve the system of linear equations below. x + y = 4 2x + 3y = 0

OpenStudy (anonymous):

x+y= 4 x=4-y add -y to both sides,cancel out y on left now plug in the value of x in the 2nd statement 2(4-y) + 3v = 0 8-2y + 3y = 0 8 +1y = 0 or 8 +y =0 y = -8 add -8 to both sides and cancel out 8 on left

OpenStudy (anonymous):

im kinda confussed.